I am trying to figure out how to elimination problems....
Not having much luck....can someone help me, point me in the right direction?
6x+2y-4=0 and 10x+7y=-8

Found 2 solutions by Alan3354, checkley77:
Answer by Alan3354(69443)

6x+2y-4=0 --> 6x+2y = 4 --> 3x+y = 2
---------
3x + y = 2 eqn 1
10x+7y = -8 eqn 2
----------
Get the same coefficient for one of the variables. Since y's coeff is 1 in eqn 1, do it.
------------
21x + 7y = 14 eqn 1 times 7
10x + 7y = -8
-------------- Subtract
11x = 22
x = 2
--------
Sub for x into eqn 1
3*2 + y = 14
y = 8
------

Answer by checkley77(12844)

6x+2y-4=0
6x+2y=4 multiply this equation by -7
10x+7y=-8 multiply this equation by 2 & add.
-42x-14y=-28
20x+14y=-16
---------------
-22x=-44
x=-44/-22
x=2 ans.
6*2+2y-4=0
12+2y-4=0
2y=4-12
2y=-8
y=-4 ans.
Proof:
10*2+7*-4=-8
20-28=-8
-8=-8